on a day when air temperature is 34°c (normal skin temperature) a cyclist is maintaining a speed of 15 km/hr. how many grams of water must this cyclist evaporate each minute to get rid of the body heat produced by his activity?


1
Expert's answer
2021-05-16T18:00:12-0400

"Fv=(1-\eta)P,"

"Mgv=\frac{(1-\eta)(Lm+cm(100°-t°))}{t},"

"m=\frac{Mgvt}{(1-\eta)(L+c(100°-t°))},"

"m=\frac{82\cdot 9.8\cdot 15/3.6\cdot 60}{(1-0.2)\cdot(2260000+4200\cdot (100-34))}=100~g."